OpenStudy (osanseviero):

Derivatives question:

OpenStudy (osanseviero):

I have a baloon of He. It has a r=2m. Then r increases by 1cm. How is the change in volume?

OpenStudy (tkhunny):

What's stopping you from simply comparing the results? \(\dfrac{4}{3}\pi\;2^{3}\;vs.\;\dfrac{4}{3}\pi\;3^{3}\)

OpenStudy (whpalmer4):

I was wondering the same...

OpenStudy (whpalmer4):

except the second radius won't be 3m...

OpenStudy (osanseviero):

Oh...I was going to do derivatives...ok..got 40 as answer, thanks :)

OpenStudy (osanseviero):

0.4

OpenStudy (osanseviero):

As a new question, it ask to do the method of diferentiation

OpenStudy (osanseviero):

So now I take the derivative and get a different answer...by a little...why?

OpenStudy (anonymous):

differentials give a linear approximation of the change

OpenStudy (whpalmer4):

what did you get for the second answer?